# Risk-Averse Learning with Delayed feedback

### v1

@@ Abstract
In real-world scenarios, the impacts of decisions may not manifest immediately. Taking these delays into account facilitates accurate assessment and management of risk in real-world environments, thereby ensuring the efficacy of strategies. In this paper, we investigate risk-averse learning using Conditional Value at Risk (CVaR) as risk measure, while incorporating delayed feedback with unknown but bounded delays. We develop two risk-averse learning algorithms that rely on one-point and two-point zeroth-order optimization approaches, respectively. The regret achieved by the algorithms is analyzed in terms of the cumulative delay and the number of total samplings. The results suggest that the two-point risk-averse learning achieves a smaller regret bound than the one-point algorithm. Furthermore, the one-point risk-averse learning algorithm attains sublinear regret under certain delay conditions, and the two-point risk-averse learning algorithm achieves sublinear regret with minimal restrictions on the delay. We provide numerical experiments on a dynamic pricing problem to demonstrate the performance of the proposed algorithms.
keywords:
online convex optimization, risk-averse, delay.
thanks: Corresponding author: ***.
@@ 1 Introduction
Risk-averse online learning addresses uncertainty and potential adverse outcomes in decision-making processes [1, 2]. Unlike traditional learning that focuses on optimizing expected outcomes, risk-averse learning minimizes potential risks. Its application includes financial investment [3], power grid management [4], and robotics [5]. The performance of online learning is usually measured using the concept of regret, which is defined as the performance loss under the algorithm relative to the best-fixed decision in hindsight. The objective is to develop algorithms that achieve sublinear regret, ensuring their effectiveness over time.
@@
Among various risk measures, CVaR is usually selected as it offers a comprehensive risk assessment by considering the tail end of the loss distribution [6]. This makes CVaR particularly useful in high-stakes applications such as finance [7], healthcare [8], and autonomous systems [9]. Since the exact gradient of the CVaR function is generally unavailable, the zeroth-order optimization approach is used to estimate the CVaR gradient. The pioneering work [10], first applies the zeroth-order algorithm to online convex optimization by querying a single function value at each iteration, known as one-point zeroth-order optimization. Building on this, various approaches, including the one-point residual [11] and the multi-point zeroth-order optimization [12, 13], have been developed to enhance estimation performance. Specifically, the one-point residual zeroth-order optimization algorithm queries function value once at each iteration and estimates gradient using the residual between two consecutive points [11]. The multi-point zeroth-order algorithm queries function values under multiple perturbed actions at each iteration and estimates gradient using differences between function values queried simultaneously [12, 13].
@@
Many real-world learning systems operate under conditions of delayed feedback. For instance, in online learning platforms, the impact of an educational intervention might be observed with a time lag. In medical diagnosis, the efficacy of a treatment plan might become apparent after several weeks or months. In online recommender systems, click-through rates are aggregated and reported back periodically. Delayed feedback complicates the learning process by introducing uncertainty and delaying the availability of information for decision-making. This, in turn, affects the ability of learning algorithms to adjust and optimize over time. Over the past decade, efforts have been devoted to delayed online learning [14, 15, 16, 17, 18, 19]. To name a few, [14] investigates online learning with feedback delayed by constant steps. The paper [15] investigates bandit convex optimization and multi-armed bandit with delayed feedback, demonstrating that the regret bound achieved by the algorithm depends on the learning horizon and cumulative feedback delay. The paper [16] investigates bandit online learning in multi-agent games and develops no-regret algorithms that ensure convergence to Nash equilibrium when individual delays are tame. Moreover, works such as [18, 19] investigate multi-armed bandit learning with delayed feedback, where the delay can be independent and identically distributed (i.i.d.) random variable with a finite expectation value [18] and i.i.d. subexponential random delays [19].
@@
In this study, we investigate risk-averse learning with delayed feedback, where the delay is varying yet upper bounded. Firstly, since the exact CVaR gradient to perform the gradient descent is unavailable, we use zeroth-order optimization approaches to estimate it. We consider both the one-point and two-point zeroth-order methods, where the function values can be evaluated at one and two points, respectively. In both scenarios, we employ the sampling strategy proposed in [20] to query multiple values of the stochastic cost under the perturbed actions per iteration, enhancing the CVaR gradient estimate accuracy. We construct the empirical distribution function of the losses using the received feedback and estimate the CVaR gradient accordingly. The estimation error of CVaR can be bounded by the difference between the true and the empirical distribution functions constructed using finite samples. In the presence of delay, the learner might receive none or multiple feedback at each iteration. To deal with this issue, we map the arrival feedback from real time slots to virtual time slots and sort them according to the reception order. This helps quantify the error caused by using delayed CVaR gradient estimates. Unlike existing literature, such as [15], which assumes that all feedback arrives before the iteration horizon ends, we address the scenario where the delay might render some final packets become unavailable. We analyze the regret bound achieved by the algorithms in terms of the cumulative delay and total sampling number. The results indicate that: 1) the regret bound decreases as the total sampling number increases or cumulative delay decreases; 2) the two-point risk-averse learning algorithm outperforms the one-point algorithm by achieving a lower regret bound; 3) the one-point algorithm achieves sublinear regret when the delay stays within a specific upper bound, while the two-point algorithm demonstrates a high tolerance for delay, imposing minimal restrictions on delay while still achieving sublinear regret.
@@
[…]
@@ 2 Preliminaries and problem statements
Consider the cost function $J_{t}(x,\xi):\mathcal{X}\times\Xi\rightarrow\mathbb{R}$ for $t=1,\dots,T$, where $x\in\mathcal{X}$ denotes decision variable with $\mathcal{X}\subseteq\mathbb{R}^{d}$ being the admissible set, $\xi\subseteq\Xi$ denotes random noise and $T$ denotes the horizon. Assume that $\mathcal{X}$ contains the ball of radius $r$ centered at the origin, i.e., $r\mathbb{B}\subseteq\mathbb{R}^{d}$. Denote the diameter of the admissible set $\mathcal{X}$ as $D_{x}=\sup_{x,y\in\mathcal{X}}\|x-y\|$.
@@ 2.1 CVaR
We use CVaR as risk measure. Suppose $J_{t}(x,\xi)$ has the cumulative distribution function $F_{t}(y)=P(J_{t}(x,\xi)\leq y)$, and is bounded by $U>0$, i.e., $|J_{t}(x,\xi)|\leq U$. Given a confidence level $\alpha\in(0,1]$, the VaR value is defined as
@@
$=\mathbb{E}_{F}\left[J_{t}(x,\xi)|J_{t}(x,\xi)\geq J_{t}^{\alpha}\right].$
@@ 2.2 Delay model
Since the exact CVaR gradient is generally unavailable, we use the zeroth-order optimization approach to estimate it. At each time slot, we perturb the action $\hat{x}_{t}=x_{t}+\delta u_{t}$ and query the loss $J_{t}(\hat{x}_{t},\xi_{t})$, where $u_{t}$ is the direction vector sampled from a unit sphere $\mathbb{S}^{d}\in\mathbb{R}^{d}$ and $\delta$ is the perturbation radius, also known as the smoothing parameter. Assume that the loss $J_{t}(\hat{x}_{t},\xi_{t})$ will be observed after $d_{t}$ time slots, where $d_{t}\geq 0$ can vary from slot to slot. Denote $J_{s|t}(\hat{x}_{s|t})$ as the loss incurred at time slot $s$ under the action $\hat{x}_{s}$ and be observed at time slot $t$. At each time slot, the learner collects the losses and the perturb direction vectors, which is denoted as $\mathcal{L}_{t}=\{J_{s|t}(\hat{x}_{s|t}),u_{s|t}|s+d_{s}=t,s\leq t\}$. When the learner receives no loss at time $t$, we have $\mathcal{L}_{t}=\emptyset$. When the learner receives losses from $\Delta_{t}$ time slots, we sort the received losses according to the reception time and denote it as $J_{s_{1}|t}(\hat{x}_{s_{1}|t}),\dots,J_{s_{\Delta_{t}}|t}(\hat{x}_{s_{\Delta_{t}}|t})$. Then $\mathcal{L}_{t}=\{J_{s_{n}|t}(\hat{x}_{s_{n}|t}),u_{s_{n}|t}\ |\ s_{n}+d_{s_{n}}=t,s_{n}\leq t,n=1,\dots,\Delta_{t}\}$.
@@ 2.3 Problem statement
We provide the following assumptions.
Assumption 1.
Assume that the delay is bounded, i.e., $d_{t}\leq\bar{d}$, for $t\in\{1,\dots,T\}$, with $\bar{d}\geq 0$.
Assumption 2.
@@
We use the cumulative loss under the performed actions against the best action in hindsight to evaluate the performance of the algorithm:
$\mathrm{R}(T)=\sum_{t=1}^{T}C_{t}(\hat{x}_{t})-\sum_{t=1}^{T}C_{t}(x^{\ast}),$ (1)
where $\hat{x}_{t}$ is the action generated by the designed algorithm at time slot $t$, and $x^{\ast}=\operatorname*{arg\,min}_{x\in\mathcal{X}}\sum_{t=1}^{T}C_{t}(x)$ is the optimal single action over the horizon $T$. This paper aims to design risk-averse learning algorithms that achieve a sublinear regret, i.e., $\lim_{T\rightarrow\infty}\frac{{\rm R}(T)}{T}=0$.
@@ 3 Main result
To estimate the CVaR gradient, a smoothed approximation of the CVaR function is constructed as
@@
The gradient of the smoothed function is given by (see [10]):
@@
The following lemmas are provided for the regret bound analysis.
Lemma 1.
@@
Lemma 2.
[1] Given Assumption 3, $C_{t}^{\delta}(x)$ is $L_{0}$-Lipschitz in $x$, i.e., $|C_{t}^{\delta}(x)-C_{t}(x)|\leq\delta L_{0}$.
Lemma 3.
[20] Let $F$ and $G$ be two cumulative distribution functions of two random variables and the random variables are bounded by $U$. Then we have that
@@ 3.1 One-point risk-averse learning
Given the difficulty of estimating the CVaR gradient from only one sample, we aggregate multiple function value samples to perform the estimation. The sampling strategy [20] is given by $m_{t}=\phi(t)$, which satisfies
$\sum_{t=1}^{T}\frac{1}{\sqrt{\phi(t)}}\leq cT^{1-\frac{a}{2}},$ (5)
with the tuning parameters $a,c>0$. The total sampling number over the iteration horizon $T$ increases with $a$. At each time slot, we query the stochastic cost for $m_{t}$ times under the perturbed action and obtain $J(\hat{x}_{t},\xi^{i}),~{}i=1,\ldots,m_{t}$. We assume that the samples queried at each time step are grouped together and experience the same delay. At time slot $t$, the learner receives
$\mathcal{L}_{t}$ =$=\{J_{s_{n}|t}(\hat{x}_{s_{n}|t},\xi_{s_{n}|t}^{i}),u_{s_{n}|t}\ |\ s_{n}+d_{s_{n}}=t,$
@@
The empirical distribution function constructed using the samples is given as
@@
for $n=1,\dots,\Delta_{t}$. Then the CVaR estimate is ${\rm CVaR}_{\alpha}[\hat{F}_{s_{n}|t}]$ and the corresponding CVaR gradient estimate is
@@
for $n=1,\dots,\Delta_{t}$ and $t=1,\dots,T$. When the learner receives no feedback at time slot $t$, the risk-averse learning algorithm updates as $x_{k+1}=x_{k}$. When the learner receives multiple feedbacks, we set $x_{t}^{0}=x_{t}$, and then the gradient descent update processes as
@@
where $\mathcal{P}_{\mathcal{X}^{\delta}}(x):=\operatorname*{arg\,min}_{y\in\mathcal{X}^{\delta}}\|x-y\|^{2}$ denotes the projection operation with $\mathcal{X}^{\delta}=\{x\in\mathcal{X}|\frac{1}{1-\delta/r}x\in\mathcal{X}\}$ being the projection set. The projection keeps the actions $\hat{x}_{t}$ within the admissible set $\mathcal{X}$ as $\big{(}1-\frac{\delta}{r}\big{)}\mathcal{X}\oplus\delta\mathbb{B}=\big{(}1-\frac{\delta}{r}\big{)}\mathcal{X}\oplus\frac{\delta}{r}r\mathbb{B}\subseteq\big{(}1-\frac{\delta}{r}\big{)}\mathcal{X}\oplus\frac{\delta}{r}\mathcal{X}=\mathcal{X}$. The learner perturbs the action once at each time slot. We call it one-point risk-averse learning and summarize it in Algorithm 1.
@@
8:     collect feedback $\mathcal{L}_{t}$ as in (3.1)
@@
14:           build empirical distribution function $\hat{F}_{s_{n}|t}(y)$ as in (7)
@@
16:           estimate CVaR gradient: $\hat{g}_{t}^{n}$ as in (8)
@@
To specify the feedback to perform gradient descent at each time slot, we map the received losses to a virtual time slot according to their reception order. Namely, the $\tau$-th virtual time slot is associated with the $\tau$-th batch of arrival feedback $\{\tilde{J}_{\tau}(\tilde{x}_{\tau},\tilde{\xi}_{\tau}^{i}),i=1,\dots,m_{\tau}\}$. Denoting $\tau:=L_{t-1}+1$, for some $\tau\in\{1,\dots,T\}$, we have $\tilde{J}_{\tau}(\cdot)=J_{s_{1}|t}(\cdot),\dots,\tilde{J}_{\tau+\Delta_{t}-1}(\cdot)=J_{s_{\Delta_{t}}|t}(\cdot)$. Moreover, we use $\mu(\tau)$ to map the $\tau$-th virtual time slot to the real time slot, i.e., $\tilde{J}_{\tau}(\cdot)=J_{\mu(\tau)|\mu(\tau)+d_{\mu(\tau)}}(\cdot)=J_{\mu(\tau)}(\cdot)$. By the end of time slot $t$, the learner receives totally $L_{t}:=\sum_{s=1}^{t}\Delta_{s}$ batches of feedback. The learner receives $L_{T}$ batches of feedback over the horizon $T$. By the maximum delay assumption (Assumption 1), we have $T-\bar{d}\leq L_{T}\leq T$. To calculate the cumulative delay, we define an auxiliary variable $\tilde{s}_{\tau}=\tau-1-L_{\mu(\tau)-1}$. The following lemma is about the auxiliary variable $\tilde{s}_{\tau}$.
Lemma 4.
The following claims hold: 1) $\tilde{s}_{\tau}\geq 0$, for $\tau=\{1,\dots,L_{T}\}$; 2) $\sum_{\tau=1}^{L_{T}}\tilde{s}_{\tau}=\sum_{t=1}^{L_{T}-2\bar{d}-1}d_{t}+\bar{d}L_{T}-\frac{\bar{d}^{2}+\bar{d}}{2}$
The proof of Lemma 4 is provided in Appendix. Existing works, such as [15], generally assume that all feedback will arrive by the end of time slot $T$ or that the impact of losing final packets can be ignored when the horizon $T$ is large. This results in $\sum_{\tau=1}^{L_{T}}\tilde{s}_{\tau}=\sum_{t=1}^{t}d_{t}$. In Lemma 4, we provide the exact expression addressing the scenario where some packets may not be received by the end of the horizon.
The following theorem analyzes the regret bound achieved by Algorithm 1.
Theorem 1.
@@
    when $a<1$, selecting $\delta=T^{-\frac{a}{4}}$ and $\eta=(T+D)^{-\frac{3a}{4}}$, where $D=\sum_{t=1}^{L_{T}-2\bar{d}-1}d_{t}+\bar{d}L_{T}-\frac{\bar{d}^{2}+\bar{d}}{2}$, Algorithm 1 achieves ${\rm R}_{1}(T)=\tilde{\mathcal{O}}\Big{(}T^{1-\frac{a}{4}}+(T+D)^{\frac{3a}{4}}+(T+D)^{-\frac{3a}{4}}T^{\frac{a}{4}}D\Big{)}$ with high probability.

### v2

@@ Abstract
In real-world scenarios, risk-averse learning is valuable for mitigating potential adverse outcomes. However, the delayed feedback makes it challenging to assess and manage risk effectively. In this paper, we investigate risk-averse learning using Conditional Value at Risk (CVaR) as risk measure, while incorporating feedback with random but bounded delays. We develop two risk-averse learning algorithms that rely on one-point and two-point zeroth-order optimization approaches, respectively. The dynamic regrets of the algorithms are analyzed in terms of the cumulative delay and the number of total samplings. In the absence of delay, the regret bounds match the established bounds of zeroth-order stochastic gradient methods for risk-averse learning. Furthermore, the two-point risk-averse learning outperforms the one-point algorithm by achieving a smaller regret bound. We provide numerical experiments on a dynamic pricing problem to demonstrate the performance of the algorithms.
{IEEEkeywords}
Online convex optimization, risk-averse, dynamic regret, random delay
@@ 1 Introduction
In online learning, the decision-maker sequentially updates its decision in a changing environment relying on historical information such as observations of previous actions and costs. However, in real-world scenarios, these observations may arrive with delays and impact system performance. The performance of online learning is typically measured through regret, which quantifies the performance gap between the algorithm’s decisions and a chosen benchmark. A widely used benchmark is static regret [1], which compares against the best fixed decision in hindsight. In contrast, dynamic regret compares the algorithm’s performance against a time-varying sequence of optimal decisions, providing a more relevant measure in non-stationary or evolving settings [2, 3]. Common metrics used to analyze dynamic regret include function variation [2] and the path length of the optimal decisions [3].
@@
Unlike traditional online learning that focuses on optimizing expected outcomes, risk-averse online learning optimizes the tail of the loss distribution to mitigate catastrophic outcomes under uncertainties [4, 5]. Its application includes financial investment [6], power grid management [7], and robotics [8]. Among various risk measures, CVaR is widely adopted, as it offers a coherent risk assessment by considering the tail end of the loss distribution [9]. This makes CVaR particularly useful in high-stakes applications such as finance [10], healthcare [11], and autonomous systems [12]. Additionally, risk-averse learning often assumes access only to function evaluations, making zeroth-order optimization a natural choice for estimating the CVaR gradient. The pioneering work [1] applies zeroth-order algorithms to online convex optimization. This method evaluates the function at a single point per iteration, a procedure known as one-point zeroth-order optimization. Based on this, various approaches have been developed to enhance the estimation performance, including the one-point residual [13] and multi-point zeroth-order optimization [14, 15]. Specifically, the one-point residual zeroth-order optimization algorithm queries function value once at each iteration and estimates the gradient using the residual between two consecutive points [13]. The multi-point zeroth-order algorithm queries function values under multiple perturbed actions at each iteration and estimates the gradient based on differences between function values queried simultaneously [14, 15].
@@
Many real-world learning systems operate in environments with delayed feedback. For instance, in online learning platforms, the impact of an educational intervention might be observed with a time lag. In medical diagnosis, the efficacy of a treatment plan might become apparent after several weeks or months. In online recommender systems, click-through rates are aggregated and reported back at regular intervals. Accounting for delays is essential, as relying on outdated information degrades system performance, especially in risk-averse settings where timely data is essential for making conservative decisions. Efforts have been devoted to online learning with different delay settings [16, 17, 18, 19, 20, 21, 22, 23, 24]. To name a few, [16] investigates online learning with feedback delayed by a constant number of steps and analyzes the regret bound of the algorithm concerning the constant delay. The paper [17] investigates bandit convex optimization and multi-armed bandit with delayed feedback and analyzes the regret bound of the algorithm concerning the learning horizon and the cumulative feedback delay. The paper [18] investigates bandit online learning in multi-agent games and develops no-regret algorithms that ensure convergence to Nash equilibrium when individual delays are tame. Some work, e.g., [20, 21], investigate multi-armed bandit learning with delayed feedback, where the delays are independent and identically distributed (i.i.d.) random variables with a finite expectation [20] and subexponential [21], respectively. Notably, the above works evaluate algorithm performance using static regret. However, since delayed feedback presents greater challenges in changing environments, the study of online learning with delays in dynamic settings is more relevant. In such contexts, dynamic regret serves as a more appropriate metric. Only a few works, such as [23, 24, 22], have studied online learning under the dynamic regret framework, addressing the fixed constant delay [23, 24] and arbitrary delay [22].
@@
In this study, we investigate risk-averse learning with delayed feedback, where the delay is varying yet upper bounded. We assume access only to function evaluations and use arrival feedback to construct the empirical distribution of the stochastic costs. Based on this, we estimate the CVaR gradient using one-point and two-point zeroth-order optimization methods. The estimation error is bounded by the difference between the true and the constructed empirical distribution functions. Unlike existing literature [17], which assumes that all feedback arrives before the iteration horizon ends, we address the scenario where the delay may cause some feedback packets to be unavailable by the end of the horizon. Additionally, random delays might change the order of feedback sequence and introduce drift in the path length of the comparators. To address this issue, [22] assumes that delays are bounded and do not alter the order of feedback. In this work, we remove this In-Order assumption and provide a general regret bound under the arbitrary delay setting. To quantify the error introduced by delayed feedback in gradient descent, we sort the received feedback by their utilization order and align it along a virtual timeline. Then, we compute the cumulative delay, which is the aggregate of the delays experienced by the feedback. The dynamic regret bounds of the algorithms are analyzed in terms of the delay, the total sampling number, and the path length of comparators. By appropriately selecting the parameters, we theoretically show that: 1) the regret bound of the risk-averse learning algorithms decreases as the total sampling number increases or the delays decrease; 2) the two-point risk-averse learning algorithm outperforms the one-point algorithm by achieving a lower regret bound.
@@
Notations: Let $\|\cdot\|$ denote the $l_{2}$ norm. Let the notation $\mathcal{O}$ hide the constant and $\tilde{\mathcal{O}}$ hide constant and polylogarithmic factors of the number of iterations $T$, respectively. Let $A\oplus B=\{a+b|a\in A,b\in B\}$ denote the Minkowski sum of two sets of position vectors $A$ and $B$ in Euclidean space.
@@ 2 Preliminaries and problem statements
Consider the cost function $J_{t}(x,\xi):\mathcal{X}\times\Xi\rightarrow\mathbb{R}$ for $t=1,\dots,T$, where $x\in\mathcal{X}$ denotes the decision variable with $\mathcal{X}\subseteq\mathbb{R}^{d}$ being the admissible set, $\xi\subseteq\Xi$ denotes the random noise and $T$ denotes the horizon. Assume that $\mathcal{X}$ contains the ball of radius $r$ centered at the origin, i.e., $r\mathbb{B}\subseteq\mathbb{R}^{d}$. Denote the diameter of the admissible set $\mathcal{X}$ as $D_{x}=\sup_{x,y\in\mathcal{X}}\|x-y\|$.
@@ 2.1 CVaR
We use CVaR as the risk measure. Suppose $J_{t}(x,\xi)$ has the cumulative distribution function $F_{t}(y)=P(J_{t}(x,\xi)\leq y)$, and is bounded by $U>0$, i.e., $|J_{t}(x,\xi)|\leq U$. Given a confidence level $\alpha\in(0,1]$, the VaR value is defined as
@@
$=\mathbb{E}_{F}\left[J_{t}(x,\xi)|J_{t}(x,\xi)\geq J_{t}^{\alpha}\right].$ (1)
Since the exact CVaR gradient for learning strategies is generally unavailable, we construct its estimate using a zeroth-order optimization approach that queries the objective function at perturbed actions [1]. We perturb the action $\hat{x}=x+\delta u$ and query the loss $J_{t}(\hat{x},\xi)$, where $u$ is a vector sampled from a unit sphere $\mathbb{S}^{d}\in\mathbb{R}^{d}$ and $\delta$ is the perturbation radius, also known as the smoothing parameter.
@@ 2.2 Delay model
Assume that the loss $J_{t}(\hat{x}_{t},\xi_{t})$ will be observed after $d_{t}$ time slots, where $d_{t}\geq 0$ can vary from slot to slot. Denote $J_{s|t}(\hat{x}_{s|t})$ as the loss incurred at time slot $s$ under the action $\hat{x}_{s}$ and be observed at time slot $t$. At each time slot, the learner collects the losses and the perturb direction vectors, which is denoted as $\mathcal{L}_{t}=\{J_{s|t}(\hat{x}_{s|t}),u_{s|t}|s+d_{s}=t,s\leq t\}$. When the learner does not receive loss at time $t$, we have $\mathcal{L}_{t}=\emptyset$. When the learner receives multiple losses at time $t$, we denote $\Delta_{t}$ as the number of arrival feedback at time $t$. Furthermore, we sort the received losses according to their generating time and denote them as $J_{s_{1}|t}(\hat{x}_{s_{1}|t}),\dots,J_{s_{\Delta_{t}}|t}(\hat{x}_{s_{\Delta_{t}}|t})$, where $s_{n}$, for $n=1,\dots,\Delta_{t}$, is the generating time of the arrival feedback. Then, the collection of the learner at time $t$ is $\mathcal{L}_{t}=\{J_{s_{n}|t}(\hat{x}_{s_{n}|t}),u_{s_{n}|t}\ |\ s_{n}+d_{s_{n}}=t,s_{n}\leq t,n=1,\dots,\Delta_{t}\}$, see Fig. 1 for an illustrative example.
[…]
Figure 1: An example of delay model: the learner does not receive feedback at slot t1t-1, i.e., $\mathcal{L}_{t-1}=\emptyset$, and receives two feedbacks at slot $t$, which are generated at slots t2t-2 and $t$, respectively, i.e., $\mathcal{L}_{t}=\{J_{t-2|t},u_{t-2|t},J_{t|t},u_{t|t}\}$.
@@ 2.3 Problem statement
We consider the class of games that satisfy the following assumptions.
Assumption 1
The delay is bounded, i.e., $d_{t}\leq\bar{d}$, for $t\in\{1,\dots,T\}$, with $\bar{d}\geq 0$.
Assumption 2
@@
We evaluate the performance of the algorithm through the dynamic regret, which is defined as the cumulative loss under the performed actions against best actions in hindsight:
$\mathrm{DR}(T)=\sum_{t=1}^{T}C_{t}(\hat{x}_{t})-\sum_{t=1}^{T}C_{t}(x_{t}^{\ast}),$ (2)
where $\hat{x}_{t}$ is the action generated by the algorithm at time slot $t$, and $x_{t}^{\ast}={\textrm{argmin}}_{x\in\mathcal{X}}C_{t}(x)$ is the optimal action at time $t$. This paper aims to design risk-averse learning algorithms that achieve a sublinear regret, i.e., $\lim_{T\rightarrow\infty}{\rm DR}(T)/T=0$.
@@ 3 Main result
Since only function evaluations are available, we use the zeroth-order optimization method to estimate the CVaR gradient for the learning process. A smoothed approximation [1] of the CVaR function is constructed as
@@
Then, the gradient of the smoothed function (3) is given by
@@
The following lemmas analyze the CVaR function and the smoothed function, which are prepared for the regret bound analysis.
Lemma 1
@@
Lemma 2
[4] Given Assumption 3, $C_{t}^{\delta}(x)$ is $L_{0}$-Lipschitz continuous in $x$, i.e., $|C_{t}^{\delta}(x)-C_{t}(x)|\leq\delta L_{0}$.
Lemma 3
[25] Let $F$ and $G$ be two cumulative distribution functions of random variables, where the random variables are bounded by $U$. Then, we have that
@@ 3.1 One-point risk-averse learning
Given the difficulty of estimating the CVaR gradient with only one single sample, we sample multiple stochastic costs to perform the estimation at each iteration. Denote $m_{t}$ as the sampling number at time slot $t$. Let the sampling number $m_{t}$ satisfy
$\sum_{t=1}^{T}\frac{1}{\sqrt{m_{t}}}\leq cT^{1-\frac{a}{2}},$ (6)
where $a,c>0$ are tuning parameters. An example of the sampling policy that satisfies (6) is provided in [25], which is given as
$m_{t}=\phi(t)=\left\lceil b(\Delta_{T}-t+1)^{a}\right\rceil,$ (7)
where $b>0$ is the tuning parameter. In (6) and (7), the total sampling number over the iteration horizon $T$ increases with $a$. At each time slot, we query the stochastic cost for $m_{t}$ times under the perturbed action and obtain $J(\hat{x}_{t},\xi^{i}),~i=1,\ldots,m_{t}$. We assume that the samples queried within the same time slot are collected as one batch. At time slot $t$, the learner receives
$\mathcal{L}_{t}$ $=\{J_{s_{n}|t}(\hat{x}_{s_{n}|t},\xi_{s_{n}|t}^{i}),u_{s_{n}|t}\ |\ s_{n}+d_{s_{n}}=t,$
@@
Then, the empirical distribution function is constructed as:
@@
for $n=1,\dots,\Delta_{t}$. Accordingly, by the CVaR definition (2.1), the CVaR estimate is denoted as ${\rm CVaR}_{\alpha}[\hat{F}_{s_{n}|t}]$ and the corresponding CVaR gradient estimate is
@@
for $n=1,\dots,\Delta_{t}$ and $t=1,\dots,T$. When the learner receives multiple feedbacks, the gradient descent processes as
@@
where the initial value is $x_{t}^{0}=x_{t}$ and $\mathcal{P}_{\mathcal{X}^{\delta}}(x):={\textrm{argmin}}_{y\in\mathcal{X}^{\delta}}\|x-y\|^{2}$ denotes the projection operation with $\mathcal{X}^{\delta}=\{x\in\mathcal{X}|\frac{1}{1-\delta/r}x\in\mathcal{X}\}$ being the projection set. When the learner does not receive any feedback at time slot $t$, the risk-averse learning algorithm updates as $x_{k+1}=x_{k}$. The projection keeps the actions $\hat{x}_{t}$ within the admissible set $\mathcal{X}$ as $\big{(}1-\frac{\delta}{r}\big{)}\mathcal{X}\oplus\delta\mathbb{B}=\big{(}1-\frac{\delta}{r}\big{)}\mathcal{X}\oplus\frac{\delta}{r}r\mathbb{B}\subseteq\big{(}1-\frac{\delta}{r}\big{)}\mathcal{X}\oplus\frac{\delta}{r}\mathcal{X}=\mathcal{X}$. Since the learner perturbs the action once at each time slot, we call it one-point risk-averse learning, which is summarized in Algorithm 1.
@@
8:  Collect feedback $\mathcal{L}_{t}$, as in (3.1)
@@
14:    Build empirical distribution function $\hat{F}_{s_{n}|t}(y)$, as in (9)
@@
16:    Estimate CVaR gradient: $\hat{g}_{t}^{n}$, as in (10)
@@
To quantify the error introduced by using outdated feedback in gradient descent, it is crucial to specify the feedback utilized at each time slot. To do it, we map the received losses to virtual time slots according to their utilization order. Namely, the $\tau$-th virtual time slot is associated with the $\tau$-th batch of utilization feedback $\{\tilde{J}_{\tau}(\tilde{x}_{\tau},\tilde{\xi}_{\tau}^{i}),i=1,\dots,m_{\tau}\}$. We assume that delays may alter the arrival order of feedback. Correspondingly, as described in Section 2.2, we sort the feedback received in the same time slot by their generation times. This setting facilitates the regret analysis, as shown in Lemma 5 in the next section.
Note that $\Delta_{t}$ is the number of feedback received by the learner at time $t$, the overall number of feedback received by the learner at the end of time $t$ is $L_{t}=\sum_{k=1}^{t}\Delta_{k}$. Denoting $\tau:=L_{t-1}+1$, for some $\tau\in\{1,\dots,T\}$, we have $\tilde{J}_{\tau}(\cdot)=J_{s_{1}|t}(\cdot),\dots,\tilde{J}_{\tau+\Delta_{t}-1}(\cdot)=J_{s_{\Delta_{t}}|t}(\cdot)$. Fig. 2 provides an illustrative example of the mapping process. Moreover, we use $\mu(\tau)$ to map the $\tau$-th virtual time slot to the real time slot, i.e., $\tilde{J}_{\tau}(\cdot)=J_{\mu(\tau)|\mu(\tau)+d_{\mu(\tau)}}(\cdot)=J_{\mu(\tau)}(\cdot)$. In this study, we analyze the regret bound of the algorithm in terms of the cumulative delay and the iteration horizon. To compute the cumulative delay, we define an auxiliary variable $\tilde{s}_{\tau}=\tau-1-L_{\mu(\tau)-1}$.
In the following lemma, we calculate the cumulative delay, accounting for the possibility that some packets may not be received by the end of the horizon.
[…]
Figure 2: Mapping from the real timeline to the virtual timeline.
Lemma 4
The following claims hold: 1) $\tilde{s}_{\tau}\geq 0$, for $\tau=\{1,\dots,L_{T}\}$; 2) $\sum_{\tau=1}^{L_{T}}\tilde{s}_{\tau}=\sum_{t=1}^{L_{T}-2\bar{d}-1}d_{t}+\bar{d}L_{T}-\frac{\bar{d}^{2}+\bar{d}}{2}$.
The proof of Lemma 4 is provided in Appendix.
Remark 1
Some work, e.g., [17, 22], assume that all feedback arrives by the end of time slot $T$ or that the impact of losing final packets is negligible when the horizon $T$ is sufficiently large. This assumption yields $\sum_{\tau=1}^{L_{T}}\tilde{s}_{\tau}=\sum_{t=1}^{T}d_{t}$, which is the cumulative delay over time horizon $T$. According to Lemma 4, we define a new cumulative delay:
$D:=\sum_{t=1}^{L_{T}-2\bar{d}-1}d_{t}+\bar{d}L_{T}-\frac{\bar{d}^{2}+\bar{d}}{2},$ (12)
which accounts for the practical scenario where some feedback may not arrive before time $T$.
Before presenting the main results, we define the path length of optimal decisions, also known as the path length of comparators, which serves as a measure for dynamic regret analysis.
Definition 1
The path length of the comparison sequence is defined as $\mathcal{P}_{T}=\sum_{t=1}^{T-1}\|x_{t}^{\ast}-x_{t+1}^{\ast}\|$, where $x_{t}^{\ast}=\arg\min_{x\in\mathcal{X}^{\delta}}C_{t}(x)$, for all $t$, is the optimal decision at time slot $t$.
Given Definition 1, the following theorem analyzes the regret bound of Algorithm 1.
Theorem 1
@@
    when $a<1$, selecting $\delta=T^{-\frac{a}{4}}\sqrt{\bar{d}P_{T}}$ and $\eta=(T+D)^{-\frac{3a}{4}}\sqrt{\bar{d}P_{T}}$. Algorithm 1 achieves ${\rm DR}_{1}(T)=\tilde{\mathcal{O}}\Big{(}\big{(}T^{1-\frac{a}{4}}+(T+D)^{\frac{3a}{4}}\big{)}\sqrt{\bar{d}P_{T}}+(T+D)^{-\frac{3a}{4}}T^{\frac{a}{4}}D\Big{)}$ with high probability.